# Appearance-based Refinement for Object-Centric Motion Segmentation

### v1

@@ C.3 Supervised adaptation: an investigation
$\checkmark$        $63.2{\scriptstyle\;{\color[rgb]{1.,.4,.4}+2.3}}$        $61.5{\scriptstyle\;{\color[rgb]{1.,.4,.4}+1.5}}$
@@
$\checkmark$ $\checkmark$        $66.7{\scriptstyle\;{\color[rgb]{1.,.4,.4}+1.7}}$        $64.1{\scriptstyle\;{\color[rgb]{1.,.4,.4}+0.3}}$
@@
Appendix D Additional quantitative results
@@
${}^{{\dagger}}$ [94] un-sup. $\checkmark$ $128\times 224$ $33.3$ $37.0$ $29.6$ $35.8$ $38.4$ $33.2$
${}^{{\dagger}}$-sup. [94] un-sup. $\checkmark$ $128\times 224$ $33.0$ $39.1$ $26.9$ $39.5$ $44.9$ $34.2$
${}^{{\dagger}}$ask R-CNN [27] un-sup. $\checkmark$ $128\times 224$ $41.1$ $42.9$ $39.2$ $52.2$ $52.0$ $52.3$
${}^{{\dagger}}$ask R-CNN [27] un-sup. $\checkmark$ $\checkmark$ $128\times 224$ $38.8$ $39.3$ $38.2$ $50.4$ $49.8$ $50.9$
@@
${}^{{\dagger}}$ideoCutLER [88] un-sup. $\checkmark$ $480\times 854$ $57.0$ $59.0$ $55.1$ $57.3$ $57.4$ $57.2$
@@
${}^{{\dagger}}$orrFlow [40] semi-sup. $\checkmark$ $480\times 854$ $55.2$ $60.0$ $50.4$ $54.0$ $54.2$ $53.7$
${}^{{\dagger}}$ [46] semi-sup. $\checkmark$ $480\times 854$ $66.9$ $70.4$ $63.4$ $65.5$ $66.2$ $64.7$
${}^{{\dagger}}$ [41] semi-sup. $\checkmark$ $480\times 854$ $70.5$ $72.7$ $68.2$ $70.9$ $71.0$ $70.8$
${}^{{\dagger}}$ [30] semi-sup. $\checkmark$ $480\times 854$ $72.9$ $75.5$ $70.4$ $73.4$ $72.9$ $74.1$
${}^{{\dagger}}$ [9] semi-sup. $\checkmark$ $480\times 854$ 75.4 76.6 74.3 78.7 77.7 79.6
@@
${}^{\ddagger}$ [42] $\checkmark$ $128\times 128$ $67.8$ $62.0$-
@@
${}^{\ddagger}$-flow [92] $\checkmark$ $128\times 224$ $72.1$ $67.6$ $70.0$
@@
${}^{\ddagger}$-TTA [92] $\checkmark$ $\checkmark$ $480\times 854$ $80.9$ $72.3$ $72.7$
@@
${}^{\ddagger}$urs $\checkmark$ $\checkmark$ $128\times 224$ 81.1 $76.6$ 81.9
@@
Appendix E Additional qualitative results
@@
Figure 9: Visualisation of our predictions at different stages. Results are shown for a DAVIS17-m sequence (top) and a YTVOS18-m sequence (bottom). Our predictions consist of four stages from top to bottom: (i) flow-predicted masks; (ii) exemplar masks selected by the mask selector (iii) masks refined by the synthetically-trained mask corrector; (iv) refined masks after the self-supervised adaptation (i.e., the final outputs). In particular, the normal-$\checkmark$ symbols are colour-coded to indicate exemplar masks selected for different objects.
@@
Figure 10: Visualisation of transparency alpha masks on DAVIS17-m (left) and YTVOS18-m (right). These alpha masks, associated with independent queries, are predicted as by-products during the feature reconstruction process and are then upsampled via bilinear interpolation to recover full resolutions. As can be observed, the alpha masks segment the entire scene into semantically meaningful regions, which supports our argument that each query corresponds to a distinct object or background region.
@@
Appendix F Potential societal impacts

### v2

@@ 0.C.3 Supervised Adaptation: An Investigation
$\checkmark$        $63.2{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+2.3}}$        $61.5{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+1.5}}$
@@
$\checkmark$ $\checkmark$        $66.7{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+1.7}}$        $64.1{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+0.3}}$
@@
0.C.4 Using SAM for Further Refinement
Effect of Different Post-Processing Steps. Table 17 compares across different post-processing approaches, for example, CRF, SAM, and HQSAM. Note that our main framework is human-label-free and only adopts CRF as the post-processing step. Furthermore, we demonstrate a complementary performance boost when applying SAM or HQSAM [38] (both trained with human annotations) as additional refinement steps.
Human Anno. CRF (HQ)SAM-based post-processing YTVOS18-m $\mathcal{J}\uparrow$ DAVIS17-m $\mathcal{J}\uparrow$
$66.9$ $66.3$
$\checkmark$ $67.1$ $67.0$
$\checkmark$ SAM |||| HQSAM $70.6$ |||| $70.9$ $70.4$ |||| $71.2$
$\checkmark$ $\checkmark$ SAM |||| HQSAM $71.1$ |||| $71.1$ $70.9$ |||| $71.3$
Table 17: Effect of different post-processing steps.
Comparison with OCLR-TTA + SAM. We apply SAM-based refinement on both our method and OCLR-TTA (previous SoTA). As shown in Table 18, our method with SAM (Ours + SAM) outperforms OCLR-TTA with SAM (OCLR-TTA + SAM), particularly on the challenging YTVOS18-m dataset.
Method Human Anno. YTVOS18-m $\mathcal{J}\uparrow$ DAVIS17-m $\mathcal{J}\uparrow$
OCLR-TTA + SAM $\checkmark$ $62.3$ $69.4$
Ours + SAM $\checkmark$ $71.1$ $70.9$
Table 18: Comparison with OCLR-TTA + SAM.
Appendix 0.D Additional Quantitative Results
@@
MG [104] un-sup. $\checkmark$ $128\times 224$ $33.3$ $37.0$ $29.6$ $35.8$ $38.4$ $33.2$
MG-sup. [104] un-sup. $\checkmark$ $128\times 224$ $33.0$ $39.1$ $26.9$ $39.5$ $44.9$ $34.2$
Mask R-CNN [32] un-sup. $\checkmark$ $128\times 224$ $41.1$ $42.9$ $39.2$ $52.2$ $52.0$ $52.3$
Mask R-CNN [32] un-sup. $\checkmark$ $\checkmark$ $128\times 224$ $38.8$ $39.3$ $38.2$ $50.4$ $49.8$ $50.9$
@@
VideoCutLER [97] un-sup. $\checkmark$ $480\times 854$ $57.0$ $59.0$ $55.1$ $57.3$ $57.4$ $57.2$
@@
CorrFlow [47] semi-sup. $\checkmark$ $480\times 854$ $55.2$ $60.0$ $50.4$ $54.0$ $54.2$ $53.7$
UVC [52] semi-sup. $\checkmark$ $480\times 854$ $66.9$ $70.4$ $63.4$ $65.5$ $66.2$ $64.7$
MAST [46] semi-sup. $\checkmark$ $480\times 854$ $70.5$ $72.7$ $68.2$ $70.9$ $71.0$ $70.8$
CRW [34] semi-sup. $\checkmark$ $480\times 854$ $72.9$ $75.5$ $70.4$ $73.4$ $72.9$ $74.1$
DINO [11] semi-sup. $\checkmark$ $480\times 854$ 75.4 76.6 74.3 78.7 77.7 79.6
@@
SIMO [48] $\checkmark$ $128\times 128$ $67.8$ $62.0$-
@@
OCLR-flow [102] $\checkmark$ $128\times 224$ $72.1$ $67.6$ $70.0$
@@
OCLR-TTA [102] $\checkmark$ $\checkmark$ $480\times 854$ $80.9$ $72.3$ $72.7$
@@
Ours $\checkmark$ $\checkmark$ $128\times 224$ 81.1 $76.6$ 81.9
@@
Appendix 0.E Additional Qualitative Results
@@
Figure 9: Visualisation of our predictions at different stages. Results are shown for a DAVIS17-m sequence (top) and a YTVOS18-m sequence (bottom). Our predictions consist of four stages from top to bottom: (i) flow-predicted masks; (ii) exemplar masks selected by the mask selector (iii) masks refined by the synthetically-trained mask corrector; (iv) refined masks after the self-supervised adaptation (i.e., the final outputs). In particular, the $\checkmark$ symbols are colour-coded to indicate exemplar masks selected for different objects.
@@
Figure 10: Visualisation of transparency alpha masks on DAVIS17-m (left) and YTVOS18-m (right). These alpha masks, associated with independent queries, are predicted as by-products during the feature reconstruction process and are then upsampled via bilinear interpolation to recover full resolutions. As can be observed, alpha masks segment the entire scene into semantically meaningful regions, which supports our argument that each query corresponds to a distinct object or background region.
@@
Appendix 0.F Potential Societal Impacts